OpenStudy (aaronandyson):

the resultant between 2P and sqrt 2P is sqrt 10 P. What is the angle between them? @michele_liano

OpenStudy (aaronandyson):

@Michele_Laino

OpenStudy (michele_laino):

I think that we have to solve this equation: \[{\left( {2P} \right)^2} + 2P + 2\left( {2P} \right)\sqrt {2P} \cos \theta = 10P\]

OpenStudy (michele_laino):

please wait I have made a typo

OpenStudy (michele_laino):

here it is the right equation: \[{\left( {2P} \right)^2} + 2{P^2} + 2\left( {2P} \right)\sqrt 2 P\cos \theta = 10{P^2}\]

OpenStudy (aaronandyson):

hows there only one sqrt 2 P in that equation?

OpenStudy (michele_laino):

since the magnitudes of the involved forces, are: \(2P\), \(\sqrt 2 P\), and \(\sqrt{10} P\)

OpenStudy (aaronandyson):

sqrt 10 P = 6P^2 + 4Psqrt2Pcostheta?

OpenStudy (michele_laino):

sorry, we have this equation: \[{\left( {2P} \right)^2} + 2{P^2} - 2\left( {2P} \right)\sqrt 2 P\cos \theta = 10{P^2}\]

OpenStudy (michele_laino):

wherein I have applied the Carnot Theorem

OpenStudy (aaronandyson):

so,4P^2 = 4sqrt2P^2costheta?

OpenStudy (michele_laino):

yes! More precisely, we have: \[ - 4\sqrt 2 \;{P^2}\cos \theta = 4{P^2}\]

OpenStudy (aaronandyson):

how minus?

OpenStudy (michele_laino):

sorry I have made a typo: \[4\sqrt 2 \;{P^2}\cos \theta = 4{P^2}\]

OpenStudy (michele_laino):

if I divide both sides by \(4 \sqrt 2 P^2\), I get: \[\cos \theta = \frac{{4{P^2}}}{{4\sqrt 2 \;{P^2}}} = ...?\] please simplify

OpenStudy (aaronandyson):

1/sqrt2

OpenStudy (michele_laino):

correct! so, \[\theta = \arccos \left( {\frac{1}{{\sqrt 2 }}} \right) = ...?\]

OpenStudy (aaronandyson):

30??/

OpenStudy (michele_laino):

If we look at the tables, or we use a calculator, we find \(\theta= \pi/4\)

OpenStudy (michele_laino):

or \(\theta=45\) degrees

OpenStudy (aaronandyson):

Thanks ,sir.

OpenStudy (michele_laino):

:)